# Generalizable Prompt Learning of CLIP: A Brief Overview

### v2

@@ 1 Introduction
Large-scale pre-trained models, such as BERT [40] and T5 [78], have made significant progress in the fields of natural language processing and computer vision. Large-scale pre-training models in the field of computer vision, such as ViT [22], treat images as sequential data and learn a large amount of general visual knowledge through pre-training on large-scale image datasets. These advancements have laid the foundation for the emergence of multimodal pre-training models for image segmentation [63, 64, 7, 122, 21], 3D detection [66, 32, 15, 23], visual question answering [107, 112, 111, 109, 108, 106], and other applications [58, 99, 85, 57, 94, 56, 51].
@@ 2 Generalizable Prompt Learning
On this basis, researchers began to study the use of continuous textual prompts to replace manually constructed discrete prompts [52, 70, 5, 70] in the textual branch. Continuous prompts, also known as learnable prompts, consist of a sequence of learnable vectors. The opposite of discrete prompts is continuous prompts, which do not rely on manually designed templates or prompt words, but instead train and optimize learnable vectors on small sample data from downstream tasks. This enables the model to automatically generate the most suitable prompts for the current task during the learning process, thereby improving the adaptability and performance of the model [59, 47]. For example, prefix tuning adds task-specific learnable vector sequences before input, optimizing the learnable vector sequences on task-specific few sample data to avoid manual design prompts. This method has achieved better results in multiple downstream tasks, demonstrating the potential of continuous prompting in improving model adaptability and performance.
@@ 8 Conclusion
This article offers a concise summary of research on few-shot prompt learning based on CLIP, encompassing experimental findings and technical attributes of various methods. The primary aim of this review is to serve as a resource for newcomers in the research field focusing on generalizable prompting of CLIP across 15 datasets. Additionally, it aims to streamline the integration of this domain by researchers from diverse fields, providing insights into the methodologies and outcomes related to prompt learning with CLIP. This overview intends to aid researchers in understanding the advancements in leveraging CLIP for prompt learning.

### v5

@@ 1 Introduction
Large-scale pre-trained models, such as BERT [45] and T5 [93], have made significant progress in the fields of natural language processing and computer vision. Large-scale pre-training models in the field of computer vision, such as ViT [25], treat images as sequential data and learn a large amount of general visual knowledge through pre-training on large-scale image datasets. These advancements have laid the foundation for the emergence of multimodal pre-training models for image segmentation [77, 78, 7, 144, 24], 3D detection [80, 35, 15, 26], visual question answering [127, 132, 131, 129, 128, 126], and other applications [69, 118, 100, 68, 110, 67, 61, 116].
@@ 2 Generalizable Prompt Learning
On this basis, researchers began to study the use of continuous textual prompts to replace manually constructed discrete prompts [63, 84, 5, 84] in the textual branch. Continuous prompts, also known as learnable prompts, consist of a sequence of learnable vectors. The opposite of discrete prompts is continuous prompts, which do not rely on manually designed templates or prompt words, but instead train and optimize learnable vectors on small sample data from downstream tasks. This enables the model to automatically generate the most suitable prompts for the current task during the learning process, thereby improving the adaptability and performance of the model [73, 52]. For example, prefix tuning adds task-specific learnable vector sequences before input, optimizing the learnable vector sequences on task-specific few sample data to avoid manual design prompts. This method has achieved better results in multiple downstream tasks [17, 18, 19], demonstrating the potential of continuous prompting in improving model adaptability and performance.
@@ 4 Experimental Dataset
Dataset Classes Train Val Test
ImageNet [23] 1,000 1.28 M N/A 50,000
Caltech101 [32] 100 4,128 1,649 2,465
OxfordPets [88] 37 2,944 736 3,669
Flowers102 [85] 102 4,093 1,633 2,463
StanfordCars [49] 196 6,509 1,635 8,041
Food101 [4] 101 50,500 20,200 30,300
FGVCAircraft [82] 100 3,334 3,333 3,333
SUN397 [115] 397 15,880 3,970 19,850
EuroSAT [36] 10 13,500 5,400 8,100
DTD [16] 47 2,820 1,128 1,692
UCF101 [103] 101 7,639 1,898 3,783
ImageNet-V2 [96] 1,000 N/A N/A 10,000
ImageNet-Sketch [108] 1,000 N/A N/A 50,889
ImageNet-A [38] 200 N/A N/A 7,500
ImageNet-R [37] 200 N/A N/A 30,000
Table 4: Training and testing datasets. This dataset consists of $1$ image classification datasets and 4 variant datasets of ImageNet.
@@ 8 Conclusion
This article offers a concise summary of research on few-shot prompt learning based on CLIP, encompassing experimental findings and technical attributes of various methods. The primary aim of this review is to serve as a resource for newcomers in the research field focusing on generalizable prompting of CLIP across 15 datasets. Additionally, it aims to streamline the integration of this domain by researchers from diverse fields, providing insights into the methodologies and outcomes related to prompt learning with CLIP. This overview intends to aid researchers in understanding the advancements in leveraging CLIP for prompt learning. Further, we plan to investigate the potential of prompt learning in other tasks [104, 120, 58, 133, 86, 44, 55, 70, 112] and scenarios [62, 137, 111, 54, 56, 42, 72, 71, 60].